# Equivariant Evidential Deep Learning for Interatomic Potentials

### v1

@@ Abstract
Uncertainty quantification (UQ) is critical for assessing the reliability of machine learning interatomic potentials (MLIPs) in molecular dynamics (MD) simulations, identifying extrapolation regimes and enabling uncertainty-aware workflows such as active learning for training dataset construction. Existing UQ approaches for MLIPs are often limited by high computational cost or suboptimal performance. Evidential deep learning (EDL) provides a theoretically grounded single-model alternative that determines both aleatoric and epistemic uncertainty in a single forward pass. However, extending evidential formulations from scalar targets to vector-valued quantities such as atomic forces introduces substantial challenges, particularly in maintaining statistical self-consistency under rotational transformations. To address this, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a backbone-agnostic framework that models atomic forces and their uncertainty jointly by representing uncertainty as a full $3\times 3$ symmetric positive definite covariance tensor that transforms equivariantly under rotations. Experiments on diverse molecular benchmarks show that $\text{e}^{2}$IP provides a stronger accuracy-efficiency-reliability balance than the non-equivariant evidential baseline and the widely used ensemble method. It also achieves better data efficiency through the fully equivariant architecture while retaining single-model inference efficiency.
@@ 1 Introduction
Despite these advances, deploying MLIPs reliably in long-timescale MD remains challenging (Perez et al., 2025). Even small force errors can accumulate over millions of integration steps, leading to unphysical trajectories and unreliable conclusions (Fu et al., ). Reliable and computationally efficient uncertainty quantification (UQ) is therefore essential for safe deployment (Bilbrey et al., 2025), as well as for uncertainty-aware workflows such as data selection and adaptive model refinement.
@@
Ensembles are widely used and often regarded as a robust “gold standard” for UQ in MLIPs (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Wollschläger et al., 2023; Willow et al., 2025), but their training and inference costs scale linearly with the number of models, making them expensive for large-scale simulation. Moreover, recent work has raised concerns about the calibration reliability of ensembles in certain settings (Rahaman and others, 2021; Abe et al., 2022; Dern et al., 2024). These limitations motivate single-model UQ approaches that provide uncertainty estimates in a single forward pass. Evidential deep learning (EDL) is a promising direction: by predicting parameters of a prior distribution, it enables direct estimation of both aleatoric and epistemic uncertainty within one pass (Sensoy et al., 2018; Amini et al., 2020; Deng et al., 2023). However, most existing EDL methods target scalar properties (Soleimany et al., 2021). Xu et al. proposed an evidential interatomic potential (eIP) (Xu et al., 2025) for forces, but models uncertainty as independent per-axis variances in a fixed global frame (i.e., a diagonal covariance). While this can express axis-aligned anisotropy, it cannot capture cross-axis correlations (off-diagonal structure) that are often necessary in vector-valued error modeling. More critically, without a full tensor structure, uncertainty cannot transform consistently under rigid rotations, which breaks the coherence of the probabilistic scale and undermines the stability of uncertainty-based ranking decisions across orientations.
Rotation consistency is fundamental for vector-valued force prediction. Forces are equivariant vectors: under a rotation $R$, the target transforms as $\mathbf{F}\mapsto R\mathbf{F}$. When adopting a probabilistic EDL framework for uncertainty modeling, consistency requires not only an equivariant mean prediction but also a distributional structure that transforms accordingly. By basic statistical transformation rules, if a random vector satisfies $\mathbf{Y}^{\prime}=R\mathbf{Y}$, then its covariance must satisfy $\Sigma^{\prime}=R\Sigma R^{\top}$. Therefore, in EDL frameworks where uncertainty is represented by covariance tensors, equivariance of uncertainty is not an additional physical assumption but a requirement for distributional self-consistency.
Motivated by this observation, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), an equivariant evidential framework for vector-valued force prediction (Fig. 1). $\text{e}^{2}$IP represents force uncertainty as a full $3\times 3$ symmetric positive definite (SPD) covariance tensor, capturing directional structure and cross-axis correlations while remaining consistent under rotation, and retaining single forward pass efficiency. To guarantee strict positive definiteness and rotation-consistent covariance construction, we parameterize a symmetric matrix in the Lie algebra and map it to the SPD manifold via the matrix exponential. Built on a Normal–Inverse–Wishart (NIW) evidential model, $\text{e}^{2}$IP naturally decomposes uncertainty into aleatoric and epistemic components. To ensure numerical robustness in complex energy landscapes and long-timescale MD, we further introduce equivariance-preserving spectral stabilization and evidence regularization to mitigate overconfidence and prevent numerical degeneracy. Experiments on diverse molecular benchmarks demonstrate that $\text{e}^{2}$IP achieves a better balance of accuracy, efficiency, and reliability compared to baseline methods, providing a practical and physically consistent framework for uncertainty-aware molecular simulations.
@@
  • We propose an $SO(3)$-consistent full-covariance construction via a Lie-algebra parameterization and matrix exponential, together with equivariance-preserving spectral stabilization for numerically robust training.
@@
Figure 1: e2IP framework. An SE(3)-equivariant GNN predicts the force mean and NIW evidential parameters, and constructs an equivariant SPD covariance $\Sigma_{0}$ (via a symmetric-matrix Lie-algebra head and matrix exponential) to produce tensor-valued uncertainty in a single forward pass.
@@ 2 Preliminaries
We consider an atomic configuration $\mathcal{G}$ with atomic coordinates $\{\vec{x}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and atomic species $\{z_{i}\in\mathbb{Z}\}_{i=1}^{N}$. Our goal is to predict the atom-wise vector-valued property $\{\vec{y}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$, such as the force acting on atom $i$, while simultaneously estimating the uncertainty associated with each atomic prediction. Within the evidential deep learning framework, the uncertainty for each atom is represented by a $3\times 3$ covariance matrix.
Equivariant backbone features.
We assume an $SE(3)$-equivariant backbone $f_{\theta}$ (e.g., NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022)) that maps $\mathcal{G}$ to node-wise latent features organized in irreducible representation (irrep) spaces. For each atom $i$, the backbone outputs a collection of irrep-valued features $\{\mathbf{h}_{i}^{l}\}$, including scalars ($l=0$) and higher-order tensors ($l>0$). These features will later be used by task-specific heads to produce both the predictive mean and an uncertainty estimate.
@@ 2.1 Equivariance requirements for mean and covariance
Let $g=(R,\mathbf{t})\in SE(3)$, where $R$ is an orthogonal matrix with $\det(R)=1$ and $\mathbf{t}\in\mathbb{R}^{3}$. Since atomic forces are translation-invariant, physical consistency requires the predicted target to transform as a vector:
@@
The uncertainty associated with a vector prediction is represented by a symmetric positive definite (SPD) rank-2 tensor $\mathbf{\Sigma}_{i}\in\mathbb{S}^{3}_{++}$ (e.g., a covariance or scale matrix). Rotation consistency then requires
@@
In particular, predicting independent per-axis variances (a diagonal matrix in a fixed global basis) generally violates Eq. (2) under non-trivial rotations and cannot capture anisotropic correlations aligned with local geometric directions.
2.2 Irrep structure of symmetric rank-2 tensors
Any symmetric matrix $\mathbf{S}_{i}\in\mathbb{S}^{3}$ admits an orthogonal decomposition into an isotropic trace component and a traceless symmetric component:
$\mathbf{S}_{i}=\underbrace{\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}}_{0e}\;+\;\underbrace{\left(\mathbf{S}_{i}-\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}\right)}_{2e}.$ (3)
From the viewpoint of $SO(3)$ representation theory, the space of symmetric rank-2 tensors decomposes as $(0e\oplus 2e)$: the scalar ($0e$) channel corresponds to the trace, and the five-dimensional ($2e$) channel corresponds to the traceless symmetric part. This structure provides a natural, rotation-compatible parameterization for constructing $\mathbf{S}_{i}$ (and hence $\mathbf{\Sigma}_{0,i}$) from equivariant features.
@@ 3 Methodology
We propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass probabilistic framework for atom-wise vector regression with rotation-consistent uncertainty (Fig. 1). Given an input structure, an $SE(3)$-equivariant network jointly predicts the force mean and a full $3\times 3$ SPD uncertainty tensor for each atom. Uncertainty is modeled using a Normal–Inverse–Wishart (NIW) evidential formulation, yielding a multivariate Student-$t$ predictive distribution with explicit aleatoric and epistemic decomposition. The method is built on two key components: (i) an equivariant SPD covariance construction via a symmetric-matrix (Lie algebra) parameterization and matrix exponential, and (ii) an equivariance-preserving spectral stabilizer that bounds covariance scale during training, ensuring strict SPD validity, rotation consistency, and numerical stability for full-covariance learning. For completeness, we summarize all symbols and validity conditions in Appendix B.
@@ 3.1 Multivariate Evidential Learning Framework
We model the observation $\vec{y}_{i}$ with a Gaussian likelihood whose mean and covariance are latent:
@@
Following multivariate evidential regression (Meinert and Lavin, 2021), we place a Normal–Inverse–Wishart (NIW) prior over $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$:
@@
Physically interpretable scale parameterization.
To obtain an interpretable scale matrix from the network, we reparameterize
@@
where $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ is an SPD matrix and describes the shape of the aleatoric noise (up to a scalar factor controlled by $\nu_{i}$).
@@
A full derivation and the expanded negative log-likelihood are given in Appendix C.
@@
$\mathbf{U}_{\text{ale}}$ reflects observation noise, while $\mathbf{U}_{\text{epi}}$ captures uncertainty in the mean induced by limited data, and both terms transform equivariantly under rotations when $\mathbf{\Sigma}_{0,i}$ does.
@@
3.2 Equivariant SPD covariance via Lie algebra
A central challenge is predicting $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ while preserving rotation equivariance ($\mathbf{\Sigma}_{0,i}\mapsto R\mathbf{\Sigma}_{0,i}R^{\top}$). Standard parameterizations (e.g., Cholesky) can be implementation-dependent and may break equivariance.
Lie-algebra to SPD via matrix exponential.
We predict coefficients in the irreducible representation space $(0e\oplus 2e)$, map them to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$, and define
@@ 3.2 Equivariant SPD covariance via Lie algebra
Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the predicted coefficients, where $s_{i}$ is the $l=0$ channel and $\mathbf{t}_{i}$ parameterizes the $l=2$ (symmetric traceless) channel. We use a fixed linear map to obtain $\mathcal{S}_{i}\in\mathfrak{sym}(3)$; details (including the constant map $\mathbf{Q}$ and vectorization convention) are provided in Appendix D.

### v2

@@ Abstract
Uncertainty quantification (UQ) is critical for the reliable deployment of machine learning interatomic potentials (MLIPs) in molecular dynamics (MD) simulations. Evidential deep learning (EDL) offers a theoretically grounded single-model alternative to expensive ensemble methods, estimating both aleatoric and epistemic uncertainty in a single forward pass. However, extending EDL from scalar targets to vector-valued atomic forces introduces a fundamental challenge: the predicted covariance must transform equivariantly under rotations to maintain statistical self-consistency. We propose $\text{e}^{2}$IP, a backbone-agnostic evidential framework that represents force uncertainty as a full $3\times 3$ symmetric positive definite (SPD) covariance tensor, constructed via a tangent-space parameterization and Riemannian exponential map that jointly guarantees positive definiteness and $SO(3)$ equivariance. Experiments on liquid water, MD22, rMD17, and silica glass benchmarks show that $\text{e}^{2}$IP achieves better calibration (CE, ES, NLL) than deep ensembles — including heteroskedastic variants — while retaining single-model inference efficiency with ${\sim}4.5\times$ speedup.
@@ 1 Introduction
Despite these advances, deploying MLIPs reliably in long-timescale MD remains challenging (Perez et al., 2025). Force errors can become problematic when simulations enter out-of-distribution configurations, where ML force fields may produce pathological energy or force predictions (Fu et al., ). Reliable and computationally efficient uncertainty quantification (UQ) is therefore essential for safe deployment (Bilbrey et al., 2025), as well as for uncertainty-aware workflows such as data selection and adaptive model refinement.
@@
Ensembles are widely used and often regarded as a robust “gold standard” for UQ in MLIPs (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Wollschläger et al., 2023; Willow et al., 2025), but their training and inference costs scale linearly with the number of models, making them expensive for large-scale simulation. Moreover, recent work has suggested that scaling a single model can be equally effective as full ensembling in certain settings, and that ensemble calibration may require careful post-hoc adjustment (Rahaman and others, 2021; Abe et al., 2022; Dern et al., 2024). These limitations motivate single-model UQ approaches. Evidential deep learning (EDL) is a promising direction: by predicting parameters of a prior distribution, it estimates both aleatoric and epistemic uncertainty in one forward pass (Sensoy et al., 2018; Amini et al., 2020; Deng et al., 2023). However, most EDL methods target scalar properties (Soleimany et al., 2021). Xu et al. proposed eIP (Xu et al., 2025) for forces, but model uncertainty as independent per-axis variances (diagonal covariance in a fixed global frame). This cannot capture cross-axis correlations and, more critically, cannot transform consistently under rigid rotations — breaking the coherence of the probabilistic model across orientations.
This limitation is not merely aesthetic. Forces are equivariant vectors ($\mathbf{F}\mapsto R\mathbf{F}$), so by standard transformation rules the covariance must satisfy $\Sigma^{\prime}=R\Sigma R^{\top}$. Equivariance of the uncertainty tensor is therefore not an additional physical assumption but a requirement for distributional self-consistency.
We propose $\text{e}^{2}$IP (Fig. 1), an equivariant evidential framework that represents force uncertainty as a full $3\times 3$ SPD covariance tensor. Positive definiteness and $SO(3)$ equivariance are jointly guaranteed by parameterizing a symmetric matrix in the tangent space of $\mathbb{S}_{++}^{3}$ and mapping it via the Riemannian exponential (Moakher, 2005; Arsigny et al., 2007). Built on a Normal–Inverse–Wishart (NIW) evidential model, $\text{e}^{2}$IP decomposes uncertainty into aleatoric and epistemic components in a single forward pass. An equivariance-preserving spectral stabilizer prevents the numerical degeneracy that otherwise causes training to diverge. Experiments on four molecular benchmarks show that $\text{e}^{2}$IP achieves better calibration and ranking quality than deep ensembles — including heteroskedastic (MVE) and shallow variants — at ${\sim}4.5\times$ lower inference cost.
@@
  • We propose an $SO(3)$-consistent full-covariance construction via a tangent-space parameterization and Riemannian exponential map on $\mathbb{S}_{++}^{3}$, together with equivariance-preserving spectral stabilization for numerically robust training.
@@
Figure 1: e2IP framework. An SE(3)-equivariant GNN predicts the force mean and NIW evidential parameters, and constructs an equivariant SPD covariance $\Sigma_{0}$ (via a tangent-space parameterization and Riemannian exponential map) to produce tensor-valued uncertainty in a single forward pass.
@@ 2 Preliminaries
This section establishes the notation and the two key requirements—equivariance and positive definiteness—that motivate our covariance construction.
@@
We consider an atomic configuration $\mathcal{G}$ with coordinates $\{\vec{x}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and species $\{z_{i}\in\mathbb{Z}\}_{i=1}^{N}$. The goal is to predict atom-wise forces $\{\vec{y}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ together with a per-atom $3\times 3$ covariance matrix representing the uncertainty of each force prediction.
Equivariant backbone.
We assume an $SE(3)$-equivariant graph neural network $f_{\theta}$ (e.g., NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022)) that maps $\mathcal{G}$ to node-wise features organized by angular momentum order $l$: for each atom $i$, the backbone outputs features $\{\mathbf{h}_{i}^{l}\}$ including scalars ($l=0$) and higher-order tensors ($l>0$), each transforming as an irreducible representation (irrep) of $SO(3)$. Task-specific heads then map these features to the force mean and uncertainty parameters.
@@ 2.1 Equivariance requirements for mean and covariance
Let $g=(R,\mathbf{t})\in SE(3)$ with $R\in SO(3)$. Forces are equivariant vectors: rotating a molecular configuration rotates the forces accordingly, so the predicted force mean $\vec{\gamma}_{i}\in\mathbb{R}^{3}$ (the model’s best estimate of the force on atom $i$) must satisfy
@@
The same requirement extends to uncertainty. If the force uncertainty is represented by a covariance matrix $\mathbf{\Sigma}_{i}\in\mathbb{S}^{3}_{++}$ (the set of $3\times 3$ symmetric positive definite matrices), it must transform as a rank-2 tensor:
@@
Predicting independent per-axis variances (a diagonal matrix in a fixed global basis) violates Eq. (2) under non-trivial rotations and cannot capture anisotropic correlations aligned with local geometric directions. This is the central limitation of existing diagonal evidential approaches (Xu et al., 2025).
2.2 Irrep decomposition of symmetric matrices
To construct equivariant symmetric matrices from backbone features, we leverage $SO(3)$ representation theory. Any $\mathbf{S}_{i}\in\mathrm{sym}(3)$ decomposes into an isotropic (trace) and a traceless part:
$\mathbf{S}_{i}=\underbrace{\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}}_{\text{scalar }(0e)}\;+\;\underbrace{\left(\mathbf{S}_{i}-\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}\right)}_{\text{traceless }(2e)}.$ (3)
The scalar ($0e$) channel captures the isotropic scale (1 degree of freedom), and the five-dimensional traceless ($2e$) channel captures the anisotropic shape. This $(0e\oplus 2e)$ structure provides a natural, rotation-compatible parameterization for constructing $\mathbf{S}_{i}$ from equivariant features, which we exploit in Sec. 3.2.
@@ 3 Methodology
We propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass probabilistic framework for atom-wise vector regression with rotation-consistent uncertainty (Fig. 1). Given an input structure, an $SE(3)$-equivariant network jointly predicts the force mean and a full $3\times 3$ SPD uncertainty tensor for each atom. Uncertainty is modeled using a Normal–Inverse–Wishart (NIW) evidential formulation (Meinert and Lavin, 2021), yielding a multivariate Student-$t$ predictive distribution with explicit aleatoric and epistemic decomposition. The method builds on existing components — SE(3)-equivariant GNN backbones, CG tensor products for coupling irreps, and multivariate evidential regression via NIW — and introduces two novel contributions: (i) to our knowledge, the first full $3\times 3$ equivariant SPD covariance parameterization for evidential force uncertainty, achieved via a tangent-space parameterization and Riemannian exponential map that jointly guarantees positive definiteness, SO(3) equivariance, and stable NIW training; and (ii) an equivariance-preserving spectral stabilizer via coefficient-space damping that bounds eigenvalue magnitude while preserving equivariance, without which training diverges (Fig. 4).
Notation.
We set $d=3$. For atom $i$: $\vec{y}_{i}\in\mathbb{R}^{d}$ is the target force vector; $\vec{\gamma}_{i}\in\mathbb{R}^{d}$ is the predicted force mean; $\vec{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$ are latent mean and covariance (random variables under the NIW prior); $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ is the network-predicted nominal covariance (deterministic); $\nu_{i}>0$ and $\kappa_{i}>0$ are evidence parameters controlling covariance and mean uncertainty, respectively. $SE(3)$ denotes the special Euclidean group; $\mathbb{S}_{++}^{3}$ the manifold of $3\times 3$ symmetric positive definite matrices; $\mathrm{sym}(3)$ the vector space of $3\times 3$ real symmetric matrices; and “irrep” stands for irreducible representation of $SO(3)$. A comprehensive notation table is provided in Appendix B.
@@ 3.1 Multivariate Evidential Learning Framework
Having established that force uncertainty requires an equivariant SPD covariance (Sec. 2.1), we now describe the probabilistic model that jointly represents force predictions and their uncertainty, then show how it decomposes into aleatoric and epistemic components.
Generative model.
For each atom $i$, we model the observed force $\vec{y}_{i}$ as drawn from a Gaussian with unknown mean and covariance:
@@
Rather than estimating $\vec{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$ directly, we treat them as latent random variables and place a conjugate Normal–Inverse–Wishart (NIW) prior (Meinert and Lavin, 2021):
@@
Here the network predicts four quantities per atom: the force mean $\vec{\gamma}_{i}$, a nominal covariance $\mathbf{\Sigma}_{0,i}$ (a deterministic matrix that controls the shape of $\mathbf{\Sigma}_{i}$), and two scalar evidence parameters $\nu_{i}$ (covariance evidence) and $\kappa_{i}$ (mean evidence). We reparameterize the IW scale matrix as
@@
so that $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ directly describes the shape of the aleatoric noise.
@@
Marginalizing the latent $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ yields a closed-form multivariate Student-$t$ predictive distribution (derivation in Appendix C):
@@
The degrees of freedom $m_{i}$ control the heaviness of the tails: low $m_{i}$ (little evidence) produces heavy tails reflecting high uncertainty, while $m_{i}\to\infty$ recovers a Gaussian.
@@
$\mathbf{U}_{\text{ale}}$ reflects irreducible observation noise (e.g., thermal fluctuations or DFT numerical noise), while $\mathbf{U}_{\text{epi}}$ captures uncertainty in the predicted mean force due to insufficient or unrepresentative training data. Both are full $3\times 3$ SPD tensors with physical directional meaning: for example, $\mathbf{U}_{\text{ale}}$ may be anisotropic along a soft vibrational mode. Crucially, because $\mathbf{\Sigma}_{0,i}$ is constructed equivariantly (Sec. 3.2), both $\mathbf{U}_{\text{ale}}$ and $\mathbf{U}_{\text{epi}}$ transform correctly under $SO(3)$ — rotating the molecule rotates the uncertainty ellipsoid accordingly.
@@
3.2 Equivariant SPD covariance via tangent-space parameterization
The evidential framework above requires the nominal covariance $\mathbf{\Sigma}_{0,i}$ to be simultaneously (a) symmetric positive definite, (b) equivariant under $SO(3)$, and (c) parameterized in a way amenable to gradient-based optimization. Standard parameterizations such as Cholesky decomposition satisfy (a) and (c) but not (b): the triangular factor $L$ in $\mathbf{\Sigma}_{0,i}=LL^{\top}$ does not transform predictably under rotation. We address all three requirements with a tangent-space construction.
From tangent space to SPD via the Riemannian exponential map.
The manifold $\mathbb{S}_{++}^{3}$ is a Riemannian symmetric space whose tangent space at the identity is $\mathrm{sym}(3)$ (Moakher, 2005; Arsigny et al., 2007). The matrix exponential $\exp\colon\mathrm{sym}(3)\to\mathbb{S}_{++}^{3}$ is the Riemannian exponential map — a diffeomorphism that maps any unconstrained symmetric matrix to a valid SPD matrix. We predict $(0e\oplus 2e)$ irrep coefficients from the equivariant backbone, map them to a symmetric matrix $\mathcal{S}_{i}\in\mathrm{sym}(3)$ via a fixed Clebsch–Gordan basis (Sec. 2.2), and define
@@ 3.2 Equivariant SPD covariance via tangent-space parameterization
Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the predicted coefficients, where $s_{i}$ is the $l=0$ channel and $\mathbf{t}_{i}$ parameterizes the $l=2$ (symmetric traceless) channel. We use a fixed linear map to obtain $\mathcal{S}_{i}\in\mathrm{sym}(3)$; details (including the constant map $\mathbf{Q}$ and vectorization convention) are provided in Appendix D.